# Large Language Model Inference Acceleration: A Comprehensive Hardware Perspective

### v3

@@ 2.2.3. Hybrid LLM
Some other LLMs integrate attention-based and SSM-based LLMs, leveraging the complete information extraction ability of attention and the information compression capability of SSM to enhance the performance for long inputs. The Block-State-Transformer (BST) (block-state-transformer, ) integrates an SSM sublayer for long-range contextualization with a block-transformer sublayer for short-term sequence representation. This architecture combines the strengths of SSMs and block attention, and explores three distinct, fully parallelizable variants. Griffin (griffin, ) combines gated linear recurrence with local attention, featuring the Hawk layer (a type of RNN with gated linear recurrence). Jamba (jamba, ) interleaves blocks of transformer and Mamba layers, harnessing the strengths of both model families. In some of these layers, mixture of expert (MoE) is added to increase model capacity while keeping the number of active parameters manageable. Unlike BST, Infini-Transformer (infini-attention, ) combines masked local attention and long-term linear attention within a single Transformer block. This Infini-Attention mechanism incorporates compressed memory into the original attention mechanism within the constraints of limited memory and computational resources. MEGALODON (megalodon, ) is a neural architecture designed for efficient sequence modeling with infinite context length. It builds on the MEGA architecture (exponential moving average with gated attention) and introduces complex exponential moving average (CEMA), timestep normalization layer, normalized attention mechanism, and a pre-norm configuration with two-hop residuals to enhance its capability and stability.
3. Optimizations on Hardware Platforms
@@
Table 2. Existing generative LLM inference optimizations on different hardware platforms
@@
Quantization Shen et al. (shen2023efficient, ), T-MAC (wei2024t, ), Snapdragon 8 Gen3 (Snapdragon8Gen3, ), llama.cpp (llamacpp, ), NoMAD-Attention (zhang2024nomad, ) GPTQ (frantar2022gptq, ), AWQ (lin2024awq, ), SpQR (dettmers2023spqr, ), SqueezeLLM (kim2023squeezellm, ), LLM-MQ (li2023llm, ), APTQ (guan2024aptq, ), Li et al. (li2023enabling, ), LUT-GEMM (park2022lut, ), FLUTE (guo2024fast, ), FP6-LLM (xia2024fp6, ), LLM.int8 (dettmers2022gpt3, ), SmoothQuant (xiao2023smoothquant, ), QUIK (ashkboos2023towards, ), Atom (zhao2024atom, ), LLM-FP4 (liu2023llm, ) FlexRun (hur2023fast, ), HLSTransform (he2024hlstransform, ), SECDA-LLM (haris2024designing, ), Chen et al. (chen2024understanding, ), FlightLLM (zeng2024flightllm, ), EdgeLLM (huang2024edgellm, ) FIGNA (jang2024figna, ), MECLA (qin2024mecla, ), OliVe (guo2023olive, ), Li et al. (li2024quantization, ), Tender (10609625, ) Guo et al. (guo2024towards, ), TransPIM (zhou2022transpim, ), Sharda et al. (sharda2024accelerator, )
@@
Sparsity Turbo Sparse (song2024turbo, ), ProSparse (song2024prosparse, ) LLM-pruner (ma2023llm, ), SparseGPT (frantar2023sparsegpt, ), Wanda (sun2023wanda, ), E-Sparse (li2023sparse, ), Flash-LLM (xia2023flash, ), Agarwalla et al. (agarwalla2024enabling, ), DejaVu (liu2023deja, ), Sparse Transformer (child2019generating, ), Bigbird (zaheer2020big, ), StreamingLLM (xiao2023streamingllm, ), Longformer (beltagy2020longformer, ), Adaptively Sparse Attention (correia2019adaptively, ), Reformer (kitaev2020reformer, ), Sparse Flash Attention (pagliardini2023faster, ), Sparse Sinkhorn Attention (tay2020sparse, ), H2(zhang2024h2o, ) FlightLLM (zeng2024flightllm, ), EdgeLLM (huang2024edgellm, ) Spatten (wang2021spatten, ), TF-MVP (yoo2023tf, ), SOFA (wang2024sofa, ) LauWS (li2024lauws, ), HARDSEA (liu2023hardsea, ), Sharda et al. (sharda2024accelerator, )
@@
Fast Decoding LLMA (yang2023inference, ), Speculative decoding (stern2018blockwise, ), Lookahead (lookahead, ), Medusa (medusa, ), EAGLE (eagle1, ; eagle2, ), Ouroboros (zhao2024ouroboros, ), Sequoia (chen2024sequoia, ), Draft&Verify (zhang2023draft, ), Kangaroo (liu2024kangaroo, ), LayerSkip (layerskip, ), Adainfer (adainfer, ), RAEE (RAEE, ), MOD (MOD, ) C-Transformer (kim202420, ) SpecPIM (li2024specpim, )
@@
Operator Optimization FlashAttention (flashattention, ; flashattention2, ), FlashDecoding (flashdecoding, ), FlashDecoding++ (hong2024flashdecoding, ), DeepSpeed (deepspeed, ), vLLM (kwon2023vllm, ), OpenPPL (OpenPPL, ), cuBLAS (cuBLAS, ), TensorRT-LLM (tensorrt-llm, ), CUTLASS (CUTLASS, ), ByteTransformer (bytetransformer, ) LPU (moon2024lpu, ), Groq LPU (groq_lpu, ), ConSmax (liu2024consmax, ), MARCA (li2024marca, ), TCP (kim2024tcp, ), Habana Gaudi (habana_gaudi, ), Gaudi2 (habana_gaudi2, ), Gaudi3 (kaplan2024intel, ), Cerebras WSE-3 (lie2024wafer, ) PIMnast (ibrahim2024balanced, ), AttentionLego (cong2024attentionlego, ), PIM-GPT (wu2024pim, ), SAL-PIM (han2024sal, ), PipePIM (jeong2024pipepim, )
@@
Heterogeneous Cooperation Kim et al. (kim2024exploiting, ), PowerInfer (song2023powerinfer, ), PowerInfer-2 (xue2024powerinfer, ) NeuPIMs (heo2024neupims, ), IANUS (seo2024ianus, ), MoNDE (kim2024monde, ), Sharda et al. (sharda2024accelerator, ), AttAcc (choi2023unleashing, ; park2024attacc, ), Kang et al. (kang2023era, ), Kim et al. (kim2024breakthrough, ), H3D-Transformer (luo2024h3d, ), CXL-PNM (park2024lpddr, ), 3D-HI (sharma2023heterogeneous, ), SK Hynix AiMX/AiMX-xPU (AiMX, ; kim2024sk, ), Cambricon-LLM (yu2024cambricon, )
Homogeneous Cooperation He et al. (he2024distributed, ; he2024inference, ) DFX (hong2022dfx, )
3.1. Quantization
3.1.1. Overview
@@
Figure 3. Two main quantization methods: weight-only quantization and weight-activation quantization.
@@
Table 3. Quantization on CPU, GPU, FPGA, ASIC, and PIM/NDP
Hardware Weight-Only Quantization Weight-Activation Quantization
@@ 3.1.2. CPU
Weight-Only Quantization. The optimization methods of quantization on CPUs mainly focus on weight-only quantization. Shen et al. (shen2023efficient, ) leverage Intel Neural Compressor to automate the INT4 quantization process with negligible accuracy loss, supporting various quantization recipes such as GPTQ (frantar2022gptq, ), AWQ (lin2024awq, ) and TEQ (cheng2023teq, ). They further develop a tensor library tailored for CPUs, which supports mainstream instruction sets like AVX2, AVX512, AVX512_VNNI, and AMX. By providing INT4 dequantization kernels on x86 CPUs, the experimental results on mainstream LLMs including Llama2, Llama and GPT-NeoX shows the latency of token generation is ranging from 12.5 tokens/s to 50 tokens/s for models with parameters ranging from 6B to 20B, by using a single socket of 4th Generation Intel Xeon Scalable Processors (4th_Gen_Intel_Xeon, ). Qualcomm Snapdragon 8 Gen3 SoC (Snapdragon8Gen3, ) utilizes its proprietary Hexagon 700 AI processor and quantization techniques to support the efficient LLM execution. For Llama2-7B with 4-bit quantization, it achieves about 15 tokens/s. Some open-source repositories like llama.cpp (llamacpp, ) are designed for efficient LLM inference across diverse hardware platforms including CPUs, GPUs and ASICs. For Llama2-7B with 4-bit quantization, llama.cpp achieves 6 tokens/s with a single core and 32 tokens/s with eight cores on Apple M2-Ultra processors. For Llama2-7B with 2-bit quantization, llama.cpp achieves 4 tokens/s with a single core and 21 tokens/s with eight cores on M2-Ultra.
@@
Due to the overheads of weight dequantization from integer to floating, T-MAC (wei2024t, ) leverages lookup tables (LUTs) for efficient low-bit LLM inference on edge CPUs, circumventing the need for dequantization and mixed precision matrix multiplication. For Llama2-7B with 4-bit quantization, T-MAC achieves 10 tokens/s with a single core and 38 tokens/s with eight cores on Apple M2-Ultra processors (Apple_M2_Ultra, ), and 3 tokens/s on Raspberry Pi 5 (Raspberry_Pi_5, ) integrated with ARM Cortex-A76. For Llama2-7B with 2-bit quantization, T-MAC achieves 17 tokens/s with a single core and 50 tokens/s with eight cores on M2-Ultra, and 6 tokens/s on Raspberry Pi 5. Furthermore, due to the vast quantities of expensive Multiply-Add (MAD) matrix operations in the attention computations, NoMAD-Attention (zhang2024nomad, ) design an efficient attention algorithm that replaces MAD operations with in-register lookups. Through hardware-aware algorithmic designs, NoMAD-Attention achieves the computation of attention scores using repeated fast accesses to SIMD registers despite their highly limited sizes. Empirical evaluations demonstrate that for CodeLlama-7B with 4-bit quantization, NoMAD-Attention achieves 9 tokens/s with short context (e.g. 128) and 4 tokens/s with long context (e.g. 16k) on 2 Intel Xeon E5-2695 V3 14-core CPUs.
3.1.3. GPU
@@
Weight-Only Quantization. GPTQ (frantar2022gptq, ) is an one-shot weight quantization method based on approximate second-order information and error compensation, that is both highly-accurate and highly-efficient. It can quantize GPT models with 175 billion parameters in approximately four GPU hours, reducing the bitwidth down to 3-bit or 4-bit per weight, with negligible accuracy degradation relative to the uncompressed baseline. Experimental results show that the average time of per token of 3-bit OPT-175B model obtained via GPTQ running on a single A100 (80GB) is 14.1 tokens/s, which is about 3.25$\times$ faster than the FP16 version (running on 5 GPUs). On more accessible GPUs, such as the NVIDIA A6000 (48GB), the average time of per token is 7.7 tokens/s (running on 2 GPUs), which is about 4.53$\times$ faster than the FP16 version (running on 8 GPUs). AWQ (lin2024awq, ) is based on the observation that protecting 1% of salient weights whose activations are extremely large can greatly reduce quantization error. It first searches for the optimal per-channel scaling and then multiplies the salient weights with the per-channel scalings. It also reduces the bitwidth down to 3 or 4 bits per weight. Experimental results with INT4 implementation show that for Llama-2-7B, it improves the inference speed from 52 tokens/s to 194 tokens/s on RTX 4090 desktop GPU (3.73$\times$ speedup). For Llama-2-13B, the inference speed is 110 tokens/s on RTX 4090 desktop GPU. On the laptop RTX 4070 GPU (8GB), it is able to run Llama-2-13B models at 33 tokens/s, while the FP16 implementation cannot fit 7B models.
@@
To further reduce the accuracy loss for smaller models in the 1-10B parameter range, SpQR (dettmers2023spqr, ) works by identifying and isolating outlier weights, which cause particularly-large quantization errors, and storing them in higher precision like half data type (16-bit), while compressing all other weights to 3-4 bits, and achieves relative accuracy losses of less than 1% in perplexity for highly-accurate LLaMA and Falcon LLMs. Experimental results show that SpQR with 3-bit and 16-bit quantization achieves 57 tokens/s, 44 tokens/s, 22 tokens/s and 12 tokens/s on A100 GPU, respectively. Unlike SpQR, SqueezeLLM (kim2023squeezellm, ) proposes a sensitivity-based non-uniform quantization method, which searches for the optimal bit precision assignment based on second-order information. It also applies dense and sparse decomposition that stores outliers and sensitive weight values in an efficient sparse format. Experimental results show that SqueezeLLM with 3bit and 16-bit quantization achieves 63.5 tokens/s, 49.2 tokens/s, 29.1 tokens/s and 14.5 tokens/s on A6000 GPU, respectively. LLM-MQ (li2023llm, ) proposes sensitivity-based precision allocation to assign the proper bitwidth for each layer within the given budget for weight memory based on their first-order information and quantization error. It also develops an efficient CUDA core kernels to accelerate LLMs by fusing the dequantization and general matrix-vector multiplication (GEMV). LLM-MQ deploys INT4 quantized Llama2-7B model on NVIDIA T4 GPU achieves up to 1.6$\times$ end-to-end speedup compared to the pytorch FP16 baseline. APTQ (guan2024aptq, ) proposes an attention-aware 2/4-bit mixed-precision quantization for LLMs, which considers not only the second-order information of each layer’s weights, but also, for the first time, the nonlinear effect of attention outputs on the entire model. Li et al. (li2023enabling, ) are the first to propose an intra-weight mixed-precision quantization for LLMs to further reduce accuracy loss under 3-bit. By applying 2/4-bit mixed-precision quantization with memory alignment and exclusive 2-bit sparse outlier reservation with minimum speed degradation, it achieves 2.91-bit for each weight considering all scales/zeros for different models with negligible loss. Additionally, they design an asynchronous dequantization and fuse the dequantization and GEMV kernels during inference. For Llama2-7B, it achieves 45.2 tokens/s on RTX 3090 GPU and 34.0 tokens/s on RTX 2080 GPU.
@@
LUT-GEMM (park2022lut, ) proposes an efficient LUT-based GPU kernel for quantized matrix multiplication, which not only eliminates the resource-intensive dequantization process but also reduces computational costs compared to previous kernels for weight-only quantization. The impact of LUT-GEMM is facilitated by implementing high compression ratios through low-bit quantization and efficient LUT-based operations. For Llama-7B with 4-bit quantization, it achieves 163.9 tokens/s on A100 GPU, achieving a remarkable 1.64$\times$ token generation latency improvement compared to the pytorch FP16 baseline. FLUTE (guo2024fast, ) is a flexible lookup table engine for LUT-quantized LLMs, which uses offline restructuring of the quantized weight matrix to minimize bit manipulations associated with unpacking, and vectorization and duplication of the lookup table to mitigate shared memory bandwidth constraints. For Llama3-8B with 4-bit quantization, it achieves 91.3-99.8 tokens/s and 113.7-121.7 tokens/s on NVIDIA A6000 and A100 GPUs, respectively. For Llama3-8B with 3-bit quantization, it achieves 91.9-110.0 tokens/s and 117.7-135.5 tokens/s on NVIDIA A6000 and A100 GPUs, respectively.
@@
To effectively reduce the size of LLMs and preserve the model accuracy, FP6-LLM (xia2024fp6, ) proposes FP6 quantization on GPUs with TC-FPx, the first full-stack GPU kernel design scheme with unified Tensor Core support of float-point weights for various quantization bit-width. It solves the unfriendly memory access of model weights with irregular bit-width and high runtime overhead of weight de-quantization. Experimental results shows that for Llama2-13B with FP6 quantization, it achieves about 55 tokens/s on NVIDIA A100 GPU.

### v4

@@ 2.2.3 Hybrid LLM
Some other LLMs integrate attention-based and SSM-based LLMs, leveraging the complete information extraction ability of attention and the information compression capability of SSM to enhance the performance for long inputs. The Block-State-Transformer (BST) [96] integrates an SSM sublayer for long-range contextualization with a block-transformer sublayer for short-term sequence representation. This architecture combines the strengths of SSMs and block attention, and explores three distinct, fully parallelizable variants. Griffin [97] combines gated linear recurrence with local attention, featuring the Hawk layer (a type of RNN with gated linear recurrence). Jamba [98] interleaves blocks of transformer and Mamba layers, harnessing the strengths of both model families. In some of these layers, mixture of expert (MoE) is added to increase model capacity while keeping the number of active parameters manageable. Unlike BST, Infini-Transformer [99] combines masked local attention and long-term linear attention within a single Transformer block. This Infini-Attention mechanism incorporates compressed memory into the original attention mechanism within the constraints of limited memory and computational resources. MEGALODON [100] is a neural architecture designed for efficient sequence modeling with infinite context length. It builds on the MEGA architecture (exponential moving average with gated attention) and introduces complex exponential moving average (CEMA), timestep normalization layer, normalized attention mechanism, and a pre-norm configuration with two-hop residuals to enhance its capability and stability.
3 Optimizations on Hardware Platforms
@@
Table 2: Existing LLM Inference Optimizations across Different Hardware Platforms
@@
Quantization Shen et al. [101], T-MAC [102], Snapdragon 8 Gen3 [64], llama.cpp [103], NoMAD-Attention [104], Zhou et al. [105], Jayanth et al. [106], Gope et al. [107], Yu et al. [108], HeteroLLM [109], DECA [110] GPTQ [111], AWQ [112], SpQR [113], SqueezeLLM [114], LLM-MQ [115], APTQ [116], Li et al. [117], LUT-GEMM [118], FLUTE [119], FP6-LLM [120], LLM.int8 [121], SmoothQuant [122], QUIK [123], Atom [124], LLM-FP4 [125], GQSA [126], Su et al. [127], Ma et al. [128] FlexRun [129], HLSTransform [130], SECDA-LLM [131], Chen et al. [132], FlightLLM [133], EdgeLLM [134], On-Device Qwen2.5 [135], LightMamba [136], TeLLMe [137], TerEffic [138], Li et al. [139], MEADOW [140], TTD [141], LlamaF [142], LNS-LLM [143], LoopLynx [144], AccLLM [145] FIGNA [146], MECLA [147], OliVe [148], Li et al. [149], Tender [150], Cai et al. [151], FineQ [152], BitMoD [153], OFQ-LLM [154], M-ANT [155], FlexiBit [156], Anda [157], AirGun [158], Ecco [159], OwL-P [160], FGMP [161], LightRot [162] Guo et al. [163], TransPIM [164], Sharda et al. [165], SoftmAP [166], MVDRAM [167], PIM-LLM [168], ROMA [169], Lue et al. [170]
@@
Sparsity Turbo Sparse [171], ProSparse [172], SparAMX [173], DECA [110] LLM-pruner [174], SparseGPT [175], Wanda [176], E-Sparse [177], Flash-LLM [178], Agarwalla et al. [179], DejaVu [180], Sparse Transformer [181], Bigbird [182], StreamingLLM [183], Longformer [184], Adaptively Sparse Attention [185], Reformer [186], Sparse Flash Attention [187], Sparse Sinkhorn Attention [188], H2[189], shihab et al. [190], SpInfer [191], SoLA [192], R-Sparse [193] FlightLLM [133], EdgeLLM [134], ChatOPU [194], Zhang et al. [195], AccLLM [145] Spatten [196], TF-MVP [197], SOFA [198] LauWS [199], HARDSEA [200], Sharda et al. [165], FlexCiM [201]
@@
Fast Decoding Dovetail [202], ML-SpecQD [203] LLMA [204], Speculative decoding [205], Lookahead [206], Medusa [207], EAGLE [208, 209], Ouroboros [210], Sequoia [211], Draft&Verify [212], Kangaroo [213], LayerSkip [214], Adainfer [215], RAEE [216], MOD [217], SpecEE [218], AMUSD [219], PipeSpec [220], Adaptix [221], SPIN [222], PARD [223], Judge Decoding [224], Falcon [225] C-Transformer [226] SpecPIM [227]
@@
Operator Optimization V-Seek [228], FlexInfer [229] FlashAttention [230, 231], FlashDecoding [232], FlashDecoding++ [233], DeepSpeed [234], vLLM [235], OpenPPL [236], cuBLAS [237], TensorRT-LLM [238], CUTLASS [239], ByteTransformer [240], SpInfer [191], FlashFormer [241], Dong et al. [242] MEADOW [140], HAAN [243] LPU [244], Groq LPU [76], ConSmax [245], MARCA [246], TCP [247], Habana Gaudi [248], Gaudi2 [249], Gaudi3 [250], Cerebras WSE-3 [77], PICACHU [251], WaferLLM [252] PIMnast [253], AttentionLego [254], PIM-GPT [255], SAL-PIM [256], PipePIM [257], SoftmAP [166]
@@
Heterogeneous Cooperation Kim et al. [258], PowerInfer [259], PowerInfer-2 [260], Dovetail [202], Twinpilots [261], Vellaisamy et al. [262], PRIMA.CPP [263], HeteroLLM [109] TightLLM [264] GLITCHES [265], HPU [266] NeuPIMs [267], IANUS [268], MoNDE [269], Sharda et al. [165], AttAcc [270, 271], Kang et al. [272], Kim et al. [273], H3D-Transformer [274], CXL-PNM [275], 3D-HI [276], SK Hynix AiMX/AiMX-xPU [277, 278], Cambricon-LLM [279], PAISE [280], PIM-AI [281], Pyramid [282], PIM-LLM [168], Lue et al. [170], Hermes [283]
Homogeneous Cooperation He et al. [284, 285] DFX [286], LoopLynx [144] Chen et al. [287]
3.1 Quantization
3.1.1 Overview
@@
Figure 3: Two main quantization methods: weight-only quantization and weight-activation quantization.
@@
KV Cache Quantization Su et al. [127] reveals that a small subset of unusual tokens exhibit unique characteristics that deviate from the pattern that keys are distributed by channel and the Values are distributed by token, which can substantially impact quantization accuracy. So they develop a outlier token tracing method to identify these tokens accurately during the decoding process and exclude them from quantization as outlier tokens. They achieve around 480 tokens/s for Llama2-7B model on a single NVIDIA A100 GPU.
Table 3: Quantization on CPU, GPU, FPGA, ASIC, and PIM/NDP
Hardware Weight-Only Quantization Weight-Activation Quantization KV Cache Quantization
@@ 3.1.2 CPU
Weight-Only Quantization. The optimization methods of quantization on CPUs mainly focus on weight-only quantization. Shen et al. [101] leverage Intel Neural Compressor to automate the INT4 quantization process with negligible accuracy loss, supporting various quantization recipes such as GPTQ [111], AWQ [112] and TEQ [288]. They further develop a tensor library tailored for CPUs, which supports mainstream instruction sets like AVX2, AVX512, AVX512_VNNI, and AMX. By providing INT4 dequantization kernels on x86 CPUs, the experimental results on mainstream LLMs including Llama2, Llama and GPT-NeoX shows the latency of token generation is ranging from 12.5 tokens/s to 50 tokens/s for models with parameters ranging from 6B to 20B, by using a single socket of 4th Generation Intel Xeon Scalable Processors [65]. Qualcomm Snapdragon 8 Gen3 SoC [64] utilizes its proprietary Hexagon 700 AI processor and quantization techniques to support the efficient LLM execution. For Llama2-7B with 4-bit quantization, it achieves about 15 tokens/s. Some open-source repositories like llama.cpp [103] are designed for efficient LLM inference across diverse hardware platforms including CPUs, GPUs and ASICs. For Llama2-7B with 4-bit quantization, llama.cpp achieves 6 tokens/s with a single core and 32 tokens/s with eight cores on Apple M2-Ultra processors. For Llama2-7B with 2-bit quantization, llama.cpp achieves 4 tokens/s with a single core and 21 tokens/s with eight cores on M2-Ultra.
@@
Due to the overheads of weight dequantization from integer to floating, T-MAC [102] leverages lookup tables (LUTs) for efficient low-bit LLM inference on edge CPUs, circumventing the need for dequantization and mixed precision matrix multiplication. For Llama2-7B with 4-bit quantization, T-MAC achieves 10 tokens/s with a single core and 38 tokens/s with eight cores on Apple M2-Ultra processors [63], and 3 tokens/s on Raspberry Pi 5 [289] integrated with ARM Cortex-A76. For Llama2-7B with 2-bit quantization, T-MAC achieves 17 tokens/s with a single core and 50 tokens/s with eight cores on M2-Ultra, and 6 tokens/s on Raspberry Pi 5. Furthermore, due to the vast quantities of expensive Multiply-Add (MAD) matrix operations in the attention computations, NoMAD-Attention [104] design an efficient attention algorithm that replaces MAD operations with in-register lookups. Through hardware-aware algorithmic designs, NoMAD-Attention achieves the computation of attention scores using repeated fast accesses to SIMD registers despite their highly limited sizes. Empirical evaluations demonstrate that for CodeLlama-7B with 4-bit quantization, NoMAD-Attention achieves 9 tokens/s with short context (e.g. 128) and 4 tokens/s with long context (e.g. 16k) on 2 Intel Xeon E5-2695 V3 14-core CPUs.
Zhou et al. [105] apply different bitwidth to compress LLMs and deploy them on Intel Xeon CPU Max 9468. Experiments show that for OPT-13B and OPT-30B, it achieves 33 tokens/s and 15.625 tokens/s with batchsize 1. Gope et al. [107] present a group-wise non-uniform codebook-based quantization method and highly optimized kernels for ultra-low-precision quantization. Experiments show that for Llama3-8B with 4-bit quantization, it achieves 45.5 tokens/s with batchsize 1 and 184.8 tokens/s with batchsize 8 on Arm Graviton3 CPU (64 cores). Jayanth et al. [106] deploy INT4 LLM on Intel Core Ultra 7 165H Processor and achieve 3.94 tokens/s and 15.6 tokens/s for Llama3-8B and TinyLlama-1B. Yu et al. [108] deploy INT4 LLM and llama.cpp on Intel Ultra-125H and Core i9-12900K and achieve 15.625 tokens/s and 6.393 tokens/s for Llama2-7B. DECA [110] first develops an analytical performance model with a 3D visual representation that provides insights into how memory resources, vector units, and hardware matrix engines interact to deliver compressed (quantization and sparsity) GeMM. Then, DECA designs a new near-core ML-model decompression accelerator. DECA offloads tile de-sparsification and dequantization from the CPU, producing ready-to-use tiles for in-core GeMM engines. In a simulated 56-core Xeon 4 server with HBM, for Llama2-70B, DECA achieves 14-25 tokens/s with MXFP4 and BF8 with 5%-20% sparsity (bs=1) and 194-281 tokens/s (bs=16).
3.1.3 GPU
@@
Weight-Only Quantization. GPTQ [111] is an one-shot weight quantization method based on approximate second-order information and error compensation, that is both highly-accurate and highly-efficient. It can quantize GPT models with 175 billion parameters in approximately four GPU hours, reducing the bitwidth down to 3-bit or 4-bit per weight, with negligible accuracy degradation relative to the uncompressed baseline. Experimental results show that the average time of per token of 3-bit OPT-175B model obtained via GPTQ running on a single A100 (80GB) is 14.1 tokens/s, which is about 3.25$\times$ faster than the FP16 version (running on 5 GPUs). On more accessible GPUs, such as the NVIDIA A6000 (48GB), the average time of per token is 7.7 tokens/s (running on 2 GPUs), which is about 4.53$\times$ faster than the FP16 version (running on 8 GPUs). AWQ [112] is based on the observation that protecting 1% of salient weights whose activations are extremely large can greatly reduce quantization error. It first searches for the optimal per-channel scaling and then multiplies the salient weights with the per-channel scalings. It also reduces the bitwidth down to 3 or 4 bits per weight. Experimental results with INT4 implementation show that for Llama-2-7B, it improves the inference speed from 52 tokens/s to 194 tokens/s on RTX 4090 desktop GPU (3.73$\times$ speedup). For Llama-2-13B, the inference speed is 110 tokens/s on RTX 4090 desktop GPU. On the laptop RTX 4070 GPU (8GB), it is able to run Llama-2-13B models at 33 tokens/s, while the FP16 implementation cannot fit 7B models.
@@
To further reduce the accuracy loss for smaller models in the 1-10B parameter range, SpQR [113] works by identifying and isolating outlier weights, which cause particularly-large quantization errors, and storing them in higher precision like half data type (16-bit), while compressing all other weights to 3-4 bits, and achieves relative accuracy losses of less than 1% in perplexity for highly-accurate LLaMA and Falcon LLMs. Experimental results show that SpQR with 3-bit and 16-bit quantization achieves 57 tokens/s, 44 tokens/s, 22 tokens/s and 12 tokens/s on A100 GPU, respectively. Unlike SpQR, SqueezeLLM [114] proposes a sensitivity-based non-uniform quantization method, which searches for the optimal bit precision assignment based on second-order information. It also applies dense and sparse decomposition that stores outliers and sensitive weight values in an efficient sparse format. Experimental results show that SqueezeLLM with 3bit and 16-bit quantization achieves 63.5 tokens/s, 49.2 tokens/s, 29.1 tokens/s and 14.5 tokens/s on A6000 GPU, respectively. LLM-MQ [115] proposes sensitivity-based precision allocation to assign the proper bitwidth for each layer within the given budget for weight memory based on their first-order information and quantization error. It also develops an efficient CUDA core kernels to accelerate LLMs by fusing the dequantization and general matrix-vector multiplication (GEMV). LLM-MQ deploys INT4 quantized Llama2-7B model on NVIDIA T4 GPU achieves up to 1.6$\times$ end-to-end speedup compared to the pytorch FP16 baseline. APTQ [116] proposes an attention-aware 2/4-bit mixed-precision quantization for LLMs, which considers not only the second-order information of each layer’s weights, but also, for the first time, the nonlinear effect of attention outputs on the entire model. Li et al. [117] are the first to propose an intra-weight mixed-precision quantization for LLMs to further reduce accuracy loss under 3-bit. By applying 2/4-bit mixed-precision quantization with memory alignment and exclusive 2-bit sparse outlier reservation with minimum speed degradation, it achieves 2.91-bit for each weight considering all scales/zeros for different models with negligible loss. Additionally, they design an asynchronous dequantization and fuse the dequantization and GEMV kernels during inference. For Llama2-7B, it achieves 45.2 tokens/s on RTX 3090 GPU and 34.0 tokens/s on RTX 2080 GPU.
@@
LUT-GEMM [118] proposes an efficient LUT-based GPU kernel for quantized matrix multiplication, which not only eliminates the resource-intensive dequantization process but also reduces computational costs compared to previous kernels for weight-only quantization. The impact of LUT-GEMM is facilitated by implementing high compression ratios through low-bit quantization and efficient LUT-based operations. For Llama-7B with 4-bit quantization, it achieves 163.9 tokens/s on A100 GPU, achieving a remarkable 1.64$\times$ token generation latency improvement compared to the pytorch FP16 baseline. FLUTE [119] is a flexible lookup table engine for LUT-quantized LLMs, which uses offline restructuring of the quantized weight matrix to minimize bit manipulations associated with unpacking, and vectorization and duplication of the lookup table to mitigate shared memory bandwidth constraints. For Llama3-8B with 4-bit quantization, it achieves 91.3-99.8 tokens/s and 113.7-121.7 tokens/s on NVIDIA A6000 and A100 GPUs, respectively. For Llama3-8B with 3-bit quantization, it achieves 91.9-110.0 tokens/s and 117.7-135.5 tokens/s on NVIDIA A6000 and A100 GPUs, respectively.
@@
To effectively reduce the size of LLMs and preserve the model accuracy, FP6-LLM [120] proposes FP6 quantization on GPUs with TC-FPx, the first full-stack GPU kernel design scheme with unified Tensor Core support of float-point weights for various quantization bit-width. It solves the unfriendly memory access of model weights with irregular bit-width and high runtime overhead of weight de-quantization. Experimental results shows that for Llama2-13B with FP6 quantization, it achieves about 55 tokens/s on NVIDIA A100 GPU.
GQSA [126](Group Quantization and Sparse Acceleration) introduces a novel compression technique, integrating quantization and sparsification in a tightly coupled manner, leveraging GPU-friendly structured group sparsity and quantization for efficient acceleration. In the setting of INT4 and 50% sparsity, it achieves 343.43 tokens/s and 228.95 tokens/s on NVIDIA A100 for Llama2-7B and Llama2-13B models, respectively. Ma et al. [128] introduce a novel bipolar-INT data format that facilitates parallel computing and supports symmetric quantization, and implement an arbitrary precision matrix multiplication scheme that decomposes and recovers matrices at the bit level. Moreover, they design a data recovery-oriented memory management system to minimize memory access latency. For Llama2-7B, they achieve 51.94 tokens/s on a single RTX 3090 GPU.